. Find whether the discharge of capacitor through the following inductive circuit is oscillatory. C = 0.1µF,
L = 10mh, R 200 Ω. If Oscillatory, find the frequency of oscillation and resonant frequency
1
Expert's answer
2020-08-02T15:13:05-0400

As per the given question,

"c=0.1\mu F"

"L=10mh =10\times 10^{-3}h"

"R=200\Omega"

"Z=\sqrt{R^2+{X_L-X_c}}"

Resonating frequency,

"\omega=\frac{1}{\sqrt{LC}}"

Now, substituting the values,

"\omega=\frac{1}{\sqrt{0.1\times 10^{-6}\times 10\times 10^{-3}}}"

"\Rightarrow \omega =\frac{1}{\sqrt{10\times10^{-10}}}"

"\Rightarrow \omega=\frac{10^5}{\sqrt{10}}"

"\Rightarrow 2\pi f =\frac{10^5}{\sqrt{10}}"

"\Rightarrow f=\frac{10^5}{2\pi \sqrt{10}}sec^{-1}"

"\Rightarrow f=\frac{10^5}{19.87}sec^{-1}"

"\Rightarrow f=5\times 10^3 sec^{-1}"